OpenStudy (rjtrain1):

The mixed number 1 2/5 and the decimal number 1.4 can be used to represent the same value. True False

OpenStudy (sshayer):

what do you think?

OpenStudy (sshayer):

\[a \frac{ b }{ c }=\frac{ (a \times c)+b }{ c }\]

OpenStudy (rjtrain1):

false

OpenStudy (tgstudios):

lol true

OpenStudy (tgstudios):

1 2/5 = 1.4